# Learning with Calibration: Exploring Test-Time Computing of Spatio-Temporal Forecasting

### v1

@@ 5.2 Effectiveness Study (RQ1)
RMSE $5.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ ${4.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 0.40$ $5.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${5.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}}$ $\downarrow 1.18$ $4.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ ${4.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}}$ $\downarrow 1.23$ $5.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.23}}$ ${5.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}}$ $\downarrow 1.95$ 4.74±0.02 4.67±0.02 $\downarrow 1.48$ $5.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ ${5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.02}}}$ $\downarrow 1.69$
@@
MAPE(%) 27.14±0.46 26.75±0.58 $\downarrow 1.44$ $28.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ ${28.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}}$ $\downarrow 1.84$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.33}}$ ${28.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}}$ $\downarrow 2.16$ $31.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ ${31.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.01}}}$ $\downarrow 0.70$ $27.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.62}}$ ${27.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.43}}}$ $\downarrow 1.63$ $29.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}$ ${29.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.57}}}$ $\downarrow 0.91$
@@
Table 3: Performance comparison of the advanced method with ST-TTC on METR-LA benchmark.
Method MAE RMSE w/o Training Set w/o Modifying Backbone
@@
TTT-MAE gandelsman2022test 3.47±0.03 7.43±0.05
TENT wangtent 4.84±0.08 8.53±0.10
CompFormer zhang2023test 3.46±0.02 7.19±0.08
@@
OnlineTCN zinkevich2003online 4.78±0.03 8.70±0.04
FSNet pham2023learning 5.79±0.24 11.06±0.24
OneNet wen2023onenet 4.94±0.03 8.80±0.06
DOST wang2024distribution 4.38±0.02 8.26±0.03
@@
Competitive Effectiveness. We further compare our method against various advanced approaches that can leverage test-time information. Since the official source code of CompFormer and DOST is not available and uses additional data information, it leads to an unfair comparison. Nevertheless, we still include all reported METR-LA benchmark values (indicated with *) using a unified GWNet backbone, categorized into regular and online settings as presented in Table 3, based on their respective papers. Additionally, we implemented the popular TTT-MAE method as a surrogate for the unavailable TTT-ST method. Our observations are as follows: ❶ For the regular setting, our method achieves competitive results with more stable standard deviations. While other methods like CompFormer demonstrate similar performance, they often utilize more training information and computational resources. ❷ In the online setting, our method significantly outperforms existing approaches without requiring more complex model architecture modifications.
@@ 5.3 Universality Study (RQ2)
To demonstrate the universality of ST-TTC across diverse real-world scenarios, we explore various forecasting scenarios in the literature, including few-shot yuan2024spatio , long-term shao2022long , and large-scale han2024bigst .
@@
As shown in Figure 3, we present the relative performance improvement of the advanced STID model with our ST-TTC method, and give a test set prediction visualization case on the PEMS-08 dataset (see Figure 9 in the Appendix for more examples). Our observations include: ❶ ST-TTC consistently improves long-term forecasting, even more than short-term (Table LABEL:tab:rq1_consistent_performance_12), likely due to more learnable information in longer windows. ❷ As the pink and orange box shows, our method learns test-time history, capturing both the global traffic decline and local fluctuations, leading to effective calibration.
@@
Large-Scale Scenario. Beyond current regional datasets, state or national-level spatio-temporal forecasting can involve tens of thousands of stations and longer time frames. We explore large-scale scenarios using the popular LargeST benchmark (comprising SD, GBA, GLA, and CA subsets). Figure 4 illustrates the 12-step prediction performance gains of the state-of-the-art efficient spatio-temporal model PatchSTG fang2025efficient with our ST-TTC , along with a comparison of inference time complexity (For full results, see Table 7 in Appendix). We observe: ❶ Our ST-TTC consistently yields further performance improvements across all datasets, even surpassing the improvement of the second-best baseline over the PatchSTG on some datasets. ❷ The additional inference time is at most 3.82 minutes, which is a clear advantage for the achieved performance gains compared to the training time cost of up to 14 hours.
@@ 5.4 Flexibility Study (RQ3)
Figure 5: Relative improvement using our ST-TTC method in the OOD learning setting.
OOD Learning Setting. Following prior work wang2024stone , we use the SD dataset to simulate spatio-temporal shift. For the temporal dimension, we use 1-8/2019, 9-10/2019, and 11-12/2020 for training, validation, and testing, respectively. For the spatial dimension, we randomly mask 10% of nodes in the test set and consider three proportions of new nodes (10% / 15% / 20%) relative to the training node to mimic varying degrees of shift. In Figure 5, we present the 12-step average prediction performance gains of the advanced OOD learning model STONE with our ST-TTC , evaluated on all nodes and new nodes to demonstrate generalizability and scalability (Full results in Table 8). We observe that: ❶ The STONE model with ST-TTC consistently achieves performance benefits, significantly outperforming all previous settings, indicating that existing OOD models are still insufficient for true OOD generalization, while our method is highly effective. ❷ For both all and new nodes, our improvements become more pronounced as the shift increases, further demonstrating our effectiveness in handling both generalizability and scalability in challenging scenarios.
Continual Learning Setting. Following prior work chen2025eac , we used multi-period streaming spatio-temporal data to examine our ST-TTC ’s integration with continual learning method. Table 4 shows the improved 12-step forecasting of advanced continual learning models EAC and STKEC with our ST-TTC . We observed: ❶ Consistent performance gains for both models across all datasets; STKEC with ST-TTC even achieved comparable performance to best model EAC. (2) Energy-Stream achieves significant improvement over other datasets, as the ST-TTC effectively learns and calibrates temporal changes, as shown by the frequency analysis (drastic shift changes) in Figure 6.
@@
EAC $24.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ $38.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.31}}$ $31.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $14.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.17}}$ $20.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ $5.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $5.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $50.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.60}}$
@@
$23.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $37.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.27}}$ $31.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $14.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ $23.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.12}}$ $20.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$ $3.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.01}}$ $3.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.00}}$ $39.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.41}}$
$\Delta$downarrow 2.5% $\downarrow$ 1.9% $\downarrow$ 1.1% $\downarrow$ 1.4% $\downarrow$ 1.2% $\downarrow$ 1.4% $\downarrow$ 32.6% $\downarrow$ 27.8% $\downarrow$ 21.6%

### v2

@@ 5.2 Effectiveness Study (RQ1)
RMSE $5.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$ $4.98{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 0.40$ $5.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $5.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$} $\downarrow 1.18$ $4.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $4.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$} $\downarrow 1.23$ $63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.23}} 5.5$$52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$} $\downarrow 1.95$ 4.74±0.02 4.67±0.02 $\downarrow 1.48$ $5.31{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.02}}$} $\downarrow 1.69$
@@
MAPE(%) 27.14±0.46 26.75±0.58 $\downarrow 1.44$ $28.75{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $28.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$} $\downarrow 1.84$ $29.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.33}}$ $28.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$} $\downarrow 2.16$ $31.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.97}}$ $31.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.01}}$} $\downarrow 0.70$ $27.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.62}}$ $27.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.43}}$} $\downarrow 1.63$ $29.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.57}}$ $29.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.57}}$} $\downarrow 0.91$
Competitive Effectiveness. We further compare our method against various advanced approaches that can leverage test-time information. Since the official source code of CompFormer and DOST is not available and uses additional data information, it leads to an unfair comparison. Nevertheless, we still include all reported METR-LA benchmark values (indicated with *) using a unified GWNet backbone, categorized
@@
Table 3: Performance comparison of the advanced method w/ ST-TTC on METR-LA benchmark. means the results are statistically significant (t-test with p-value <0.01<0.01).
Method MAE RMSE
w/o Training
Set
w/o Modifying
Backbone
@@
TTT-MAE (Gandelsman et al., 2022) 3.47±0.03 7.43±0.05
TENT (Wang et al., 2021) 4.84±0.08 8.53±0.10
CompFormer (Zhang et al., 2023c) 3.46±0.02 7.19±0.08
@@
OnlineTCN (Zinkevich, 2003) 4.78±0.03 8.70±0.04
FSNet (Pham et al., 2023) 5.79±0.24 11.06±0.24
OneNet (Wen et al., 2023) 4.94±0.03 8.80±0.06
DOST (Wang et al., 2024c) 4.38±0.02 8.26±0.03
@@
into regular and online settings as presented in Table 3, based on their respective papers. Additionally, we implemented the popular TTT-MAE method as a surrogate for the unavailable TTT-ST method. Our observations are as follows: ❶ For the regular setting, our method achieves competitive results with more stable standard deviations. While other methods like CompFormer demonstrate similar performance, they often utilize more training information and computational resources. ❷ In the online setting, our method significantly outperforms existing approaches without requiring more complex model architecture modifications.
@@ 5.3 Universality Study (RQ2)
To demonstrate the universality of ST-TTC across diverse real-world scenarios, we explore various forecasting scenarios in the literature, including few-shot (Yuan et al., 2024c), long-term (Shao et al., 2022a), and large-scale (Han et al., 2024).
@@
As shown in Figure 3, we present the relative performance improvement of the advanced STID model with our ST-TTC method, and give a test set prediction visualization case on the PEMS-08 dataset (see Figure F.1 in the Appendix for more examples). Our observations include: ❶ ST-TTC consistently improves long-term forecasting, even more than short-term (Table LABEL:tab:rq1_consistent_performance_12), likely due to more learnable information in longer windows. ❷ As the pink and orange box shows, our method learns test-time history, capturing both the global traffic decline and local fluctuations, leading to effective calibration.
@@
Large-Scale Scenario. Beyond current regional datasets, state or national-level spatio-temporal forecasting can involve tens of thousands of stations and longer time frames. We explore large-scale scenarios using the popular LargeST benchmark (comprising SD, GBA, GLA, and CA subsets). Figure 4 illustrates the 12-step prediction performance gains of the state-of-the-art efficient spatio-temporal model PatchSTG (Fang et al., 2025) with our ST-TTC , along with a comparison of inference time complexity (For full results, see Table E.2 in Appendix). We observe: ❶ Our ST-TTC consistently yields further performance improvements across all datasets, even surpassing the improvement of the second-best baseline over the PatchSTG on some datasets. ❷ The additional inference time is at most 3.82 minutes, which is a clear advantage for the achieved performance gains compared to the training time cost of up to 14 hours.
@@ 5.4 Flexibility Study (RQ3)
Figure 5: Relative improvement using our ST-TTC in the OOD learning setting.
OOD Learning Setting. Following prior work (Wang et al., 2024a), we use the SD dataset to simulate spatio-temporal shift. For the temporal dimension, we use 1-8/2019, 9-10/2019, and 11-12/2020 for training, validation, and testing, respectively. For the spatial dimension, we randomly mask 10% of nodes in the test set and consider three proportions of new nodes (10% / 15% / 20%) relative to the training node to mimic varying degrees of shift. In Figure 5, we present the 12-step average prediction performance gains of the advanced OOD learning model STONE with our ST-TTC , evaluated on all nodes and new nodes to demonstrate generalizability and scalability (Full results in Table E.3). We observe that: ❶ The STONE model with ST-TTC consistently achieves performance benefits, significantly outperforming all previous settings, indicating that existing OOD models are still insufficient for true OOD generalization, while our method is highly effective. ❷ For both all and new nodes, our improvements become more pronounced as the shift increases, further demonstrating our effectiveness in handling both generalizability and scalability in challenging scenarios.
Continual Learning Setting. Following prior work (Chen and Liang, 2025), we used multi-period streaming spatio-temporal data to examine our ST-TTC ’s integration with continual learning method. Table 4 shows the improved 12-step forecasting of advanced continual learning models EAC and STKEC with our ST-TTC . We observed: ❶ Consistent performance gains for both models across all datasets; STKEC with ST-TTC even achieved comparable performance to best model EAC. ❷ Energy-Stream achieves significant improvement over other datasets, as the ST-TTC effectively learns and calibrates temporal changes, as shown by the frequency analysis (drastic shift changes) in Figure 6.
@@
EAC $24.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$ $38.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.31}}$ $31.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $14.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ $24.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.17}}$ $20.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $5.15{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $5.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ $50.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.60}}$
@@
$23.54{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ $37.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.27}}$ $31.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $14.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}}$ $23.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.12}}$ $20.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$ $3.47{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.01}}$ $3.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.00}}$ $39.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.41}}$
$\Delta$ $\downarrow$ 2.5% $\downarrow$ 1.9% $\downarrow$ 1.1% $\downarrow$ 1.4% $\downarrow$ 1.2% $\downarrow$ 1.4% $\downarrow$ 32.6% $\downarrow$ 27.8% $\downarrow$ 21.6%